OpenStudy (anonymous):

The diagonal of a cube is 7 sqrt 3 find the volume and its cube !! help

OpenStudy (anonymous):

check out my second answer brit

OpenStudy (anonymous):

where is it i cant find it

OpenStudy (anonymous):

here it is i assume you mean the diagonal of one side of the cube. by pythagoras (7sqrt3)^2 = 2x^2 (where x = one side of cube) 147 = 2x^2 x = sqrt(147/2) = 8.5732 vlume of cube = (8.5732)^3 = 630.13 correct to 2 dec places

OpenStudy (anonymous):

wow okay well then thank you again hope you are on tomorrow around 4 bye :)